# local non-bossiness

### v4

@@ 6. Concluding remarks
We claim that the mechanism that associates to each preference profile the school-median stable matching is locally bossy. Let $N=\{1,2,3,4\}$, $S=\{s_{1},s_{2}\}$, $q_{s_{1}}=q_{s_{2}}=2$, $\succ_{s_{1}}:2,1,3,4$, and $\succ_{s_{2}}:3,4,2,1.$ If preferences are given by
$P_{1}:2,1,\quad\quad P_{2}:2,1,\quad\quad P_{3}:1,2,\quad\quad P_{4}=1,2,$
there are three stable matchings:
@@
It follows that $\eta$ is the school-median stable matching. Now, if preferences of student $1$ change to $P^{\prime}_{1}:1,2$, the only stable matchings are $\eta$ and $\rho$. In this case, $\rho$ is the school-median stable matching. Therefore, as $\eta(1)=\rho(1)=s_{1}$ and $\eta(s_{1})\neq\rho(s_{1})$, we conclude that the school-median stable mechanism is locally bossy. $\Box$
The following diagram summarizes the causal relationships between local non-bossiness and other incentive properties:
Strategy-proofness   +   Local non-bossinessGroup strategy-proofness  Local group strategy-proofnessLocal group non-bossinessStrategy-proofnessWeak non-bossinessLocal $\bm{\times}$DA      Lemma 1Lemma $\bm{\times}$Example $\bm{\times}$Boston $\bm{\times}$Example $\bm{\times}$Example $\bm{\times}$Lemma 3   $\bm{\times}$Lemma $\bm{\times}$Lemma 3
Figure 1. On local non-bossiness and other incentive properties.
Appendix B omitted proofs
Given a student $1\in N$ and preference profiles $P,P^{\prime}\in{\mathcal{P}}$, let $\mu={\rm{DA}}(P_{1},P_{-1})$ and $\mu^{\prime}={\rm{DA}}(P^{\prime}_{1},P_{-1})$. In the following results we characterize some properties of $\mu$ and $\mu^{\prime}$.
Lemma 4.
Suppose that $\mu(1)=\mu^{\prime}(1)=\overline{s}$. If there exists $i\in N$ such that $s=\mu(i)$ and $s^{\prime}=\mu^{\prime}(i)$ are different, then $s\neq s_{0}$, $s^{\prime}\neq s_{0}$, and both schools fill their places in $\mu$ and $\mu^{\prime}$.
Proof.
Let $\widehat{\mu}$ be the matching induced by $\mu$ when student $1$ and one seat in $\overline{s}$ are eliminated (w.l.o.g., assume that $q_{s_{0}}=|N|$). Define $\widehat{\mu}^{\prime}$ analogously. Notice that $\widehat{\mu}$ and $\widehat{\mu}^{\prime}$ are stable matchings in the school choice problem $(N\setminus\{1\},S,(\succ_{r})_{r\in S},(\widehat{q}_{r})_{r\in S},P_{-1})$ where $\widehat{q}_{\overline{s}}=q_{\overline{s}}-1$ and $\widehat{q}_{r}=q_{r}$ for $r\neq\overline{s}$. Since $s=\widehat{\mu}(i)$ and $s^{\prime}=\widehat{\mu}^{\prime}(i)$ are different, the Rural Hospital Theorem (Roth, 1986) implies that $s$ and $s^{\prime}$ are schools and fill their places in both $\widehat{\mu}$ and $\widehat{\mu}^{\prime}$. [^17] In many-to-one matching problems, not only are the unassigned students the same in all stable matchings, but each school that does not fill its places enrolls the same group of students (Roth, 1986). Therefore, $s$ and $s^{\prime}$ satisfy the same properties in $\mu$ and $\mu^{\prime}$. ∎
Lemma 5.
If $\mu^{\prime}(i)P_{i}\mu(i)$, then there exists a cycle $(i_{1},\ldots.i_{r})$ that includes $i$ and satisfies:
  • (i)
    $\mu^{\prime}(i_{l})=\mu(i_{l+1})$ for all $l\in\{1,\ldots,r\}$ [modulo $r$].
  • (ii)
    $\mu^{\prime}(i_{l})P_{i_{l}}\mu(i_{l})$ for all $l\in\{1,\ldots,r\}$.
Proof.
Let $G^{*}=(V^{*},E^{*})$ be a multigraph in which $V^{*}=S$ and there is a directed edge between $s$ and $s^{\prime}$ for each student $k$ such that $\mu(k)=s$ and $\mu^{\prime}(k)=s^{\prime}$, with $s\neq s^{\prime}$. Since Lemma 4 ensures that $|\mu(s)|=q_{s}$ and $|\mu^{\prime}(s^{\prime})|=q_{s^{\prime}}$ for any $s$ and $s^{\prime}$ such that $[s,s^{\prime}]\in E^{*}$, the in-degree and the out-degree of each node in $V^{*}$ coincide. Moreover, as schools $a=\mu(i)$ and $a^{\prime}=\mu^{\prime}(i)$ are different, there is at least one edge between $a$ and $a^{\prime}$.
As a consequence, there is a cycle of schools $(s_{1},\ldots,s_{r})$ in $G^{*}$ such that $(s_{1},s_{2})=(a,a^{\prime})$. Indeed, we can form this cycle departing from node $a$ in direction to node $a^{\prime}$ and moving around the graph, because after enter a node $s_{j}\neq s_{1}$ from an edge, we are always able to exit from another edge. Since $V^{*}$ is a finite set, we will eventually return to $s_{1}$. The construction of $E^{*}$ allow us to associated to $(s_{1},\ldots,s_{r})$ a group of students $i_{1},\ldots,i_{r}$ such that $i_{1}=i$, $\mu(i_{l})=s_{l}$ for all $l\in\{1,\ldots,r\}$, and $\mu^{\prime}(i_{l})=s_{l+1}=\mu(i_{l+1})$ for all $l\in\{1,\ldots,r\}$ [modulo $r$]. In particular, although $s_{1},\ldots,s_{r}$ are not necessarily different, we can assume that the students $i_{1},\ldots,i_{r}$ are.
Since $\mu^{\prime}(i_{l})=s_{l+1}$, when $\mu^{\prime}(i_{l})P_{i_{l}}\mu(i_{l})$ the stability of $\mu$ ensures that $j\succ_{s_{l+1}}i_{l}$ for all $j\in\mu(s_{l+1})$. Hence, the stability of $\mu^{\prime}$ implies that any edge between $s_{l+1}$ and $s_{l+2}$ is associated to a student that improves we she moves from $\mu$ to $\mu^{\prime}$. In particular, $\mu^{\prime}(i_{l+1})P_{i_{l+1}}\mu(i_{l+1})$. Therefore, $\mu^{\prime}(i_{1})P_{i_{1}}\mu(i_{1})$ ensures that $\mu^{\prime}(i_{l})P_{i_{l}}\mu(i_{l})$ for all $l\in\{1,\ldots,r\}$.
We conclude that $(i_{1},\ldots,i_{r})$ satisfies properties (i) and (ii).∎
Lemma 6.
Suppose that $|N|=\sum_{s\in S}q_{s}$, $P\in{\mathcal{P}}^{*}$, and $(P^{\prime}_{1},P_{-1})\in{\mathcal{P}}^{*}$. If $\mu(1)\neq\mu^{\prime}(1)$, then ${\rm{Coll}}_{1}(P)\cap{\rm{Coll}}_{1}(P^{\prime}_{1},P_{-1})=\emptyset$.
Proof.
Since $|N|=\sum_{s\in S}q_{s}$ and everyone considers all schools admissible under the preference profiles $P$ and $(P_{1},P_{-1})$, it follows that $\overline{s}\equiv\mu(1)$ and $\hat{s}\equiv\mu^{\prime}(1)$ are two different schools. By contradiction, suppose that there exists $i_{1}\in N\setminus\{1\}$ such that $\mu(i_{1})=\overline{s}$ and $\mu^{\prime}(i_{1})=\hat{s}$. Without loss of generality, assume that $\hat{s}P_{i_{1}}\overline{s}$ (in other case, exchange the roles of $P$ and $(P^{\prime}_{1},P_{-1})$).
In the proof of Theorem 1, the fact that $\mu(1)=\mu^{\prime}(1)$ is only used to guarantee that all students who have different assignments in $\mu$ and $\mu^{\prime}$ are assigned to schools and these schools fill their quota (see Lemma 4). In the current context, although $\mu(1)\neq\mu^{\prime}(1)$, the same properties hold because DA is a stable mechanism, $|N|=\sum_{s\in S}q_{s}$, and preference profiles belong to ${\mathcal{P}}^{*}$.
Thus, after assuming the existence of $i_{1}\neq 1$ such that $\mu(i_{1})=\overline{s}$, $\mu^{\prime}(i_{1})=\hat{s}$, and $\hat{s}P_{i_{1}}\overline{s}$, the same arguments made in the proof of Theorem 1 ensure that $\mu({\overline{s}})\setminus\{1\}=\mu^{\prime}(\overline{s})\setminus\{1\}$. Since $i_{1}\in\mu({\overline{s}})\setminus\{1\}$, we conclude that $\overline{s}=\hat{s}$. A contradiction. ∎
Proof of Theorem 2. Given a priority profile $\succ$, the stability of ${\rm{DA}}^{\succ}$ ensures that it is individually rational and weakly non-wasteful. ${\rm{DA}}^{\succ}$ is also population-monotonic (Crawford, 1991), strategy-proof (Dubins and Freedman, 1981; Roth, 1982), and weakly WrARP because under deferred acceptance each school chooses the students according to its priority order. The weak local non-bossiness of ${\rm{DA}}^{\succ}$ follows from Theorem 1.
Suppose that $\Phi:{\mathcal{N}}\times\mathcal{P}\rightarrow{\mathcal{M}}^{*}$ satisfies individual rationality (IR), weak non-wastefulness (WNW), population-monotonicity (PM), strategy-proofness (SP), weak WrARP (w-WrARP), and weak local non-bossiness (WLNB).
Given $s\in S$, let $P^{s}$ be a preference relation such that $s$ is the only admissible school. Denote by $\overline{P}=(\overline{P}_{i})_{i\in{\overline{N}}}$ a preference profile such that $\overline{P}_{i}=P^{s}$ for every $i\in{\overline{N}}$. It follows from WNW that the mapping that associates to each $N\in{\mathcal{N}}$ the set $\Phi_{s}(N,\overline{P})$ is a choice function. [^18] A choice function is a mapping $C:{\mathcal{N}}\rightarrow{\mathcal{N}}$ such that, for each $N\in{\mathcal{N}}$, $C(N)$ is a non-empty subset of $N$. Moreover, WNW, PN, w-WrARP, and the Theorem 2 of Chambers and Yenmez (2018) guarantee that there exists a priority order $\succ_{s}$ defined on ${\overline{N}}$ such that the following property holds:
  • (A)
    For each $N\in{\mathcal{N}}$, if $i\in N\setminus\Phi_{s}(N,\overline{P})$, then $j\succ_{s}i$ for all $j\in\Phi_{s}(N,\overline{P})$.
Let $\succ=(\succ_{s})_{s\in S}$. We will use the property (A) to show that $\Phi={\rm{DA}}^{\succ}$.
By contradiction, assume that there is $N^{*}\in{\mathcal{N}}$ such that $\Phi(N^{*},\cdot)\neq{\rm{DA}}^{\succ}(N^{*},\cdot)$. For each $(N,P)\in{\mathcal{N}}\times{\mathcal{P}}$, let $Z(N,P)=|\{i\in N:|A(P_{i})|\leq 1\}|$ be the number of students in $N$ who find at most one school admissible. Let $P^{*}\in{\mathcal{P}}$ be such that $\Phi(N^{*},P^{*})\neq{\rm{DA}}^{\succ}(N^{*},P^{*})$ and $Z(N^{*},P^{*})\geq Z(N^{*},P)$ for all $P\in{\mathcal{P}}$ satisfying $\Phi(N^{*},P)\neq{\rm{DA}}^{\succ}(N^{*},P)$. Given $i\in N^{*}$, we claim that $\Phi_{i}(N^{*},P^{*})\neq{\rm{DA}}^{\succ}_{i}(N^{*},P^{*})$ implies that $|A(P^{*}_{i})|=1$. There are two cases to analyze:
  • Case 1: $\Phi_{i}(N^{*},P^{*})\,P^{*}_{i}\,{\rm{DA}}^{\succ}_{i}(N^{*},P^{*})$. Since $\Phi$ and ${\rm{DA}}^{\succ}$ satisfy IR, $s:=\Phi_{i}(N^{*},P^{*})$ belongs to $S$. Hence, $|A(P^{*}_{i})|\geq 1$. IR and SP imply that $\Phi_{i}(N^{*},(P^{s},P^{*}_{-i}))=s$ and ${\rm{DA}}^{\succ}_{i}(N^{*},(P^{s},P^{*}_{-i}))=s_{0}$. As a consequence, if $|A(P^{*}_{i})|>1$, we obtain that $Z(N^{*},(P^{s},P^{*}_{-i}))>Z(N^{*},P^{*})$ and $\Phi(N^{*},(P^{s},P^{*}_{-i}))\neq{\rm{DA}}^{\succ}(N^{*},(P^{s},P^{*}_{-i}))$, which contradicts the definition of $P^{*}$. Therefore, $|A(P^{*}_{i})|=1$.
  • Case 2: ${\rm{DA}}^{\succ}_{i}(N^{*},P^{*})\,P^{*}_{i}\,\Phi_{i}(N^{*},P^{*})$. The result follows from the same arguments applied in the previous case but swapping the roles of $\Phi$ and ${\rm{DA}}^{\succ}$.
Let $i\in N^{*}$ be such that $\Phi_{i}(N^{*},P^{*})\neq{\rm{DA}}^{\succ}_{i}(N^{*},P^{*})$. Since $|A(P^{*}_{i})|=1$, it follows from IR that there exists $s\in S$ such that either $s={\rm{DA}}^{\succ}_{i}(N^{*},P^{*})\,P^{*}_{i}\,\Phi_{i}(N^{*},P^{*})=s_{0}$ or $s=\Phi_{i}(N^{*},P^{*})\,P^{*}_{i}\,{\rm{DA}}^{\succ}_{i}(N^{*},P^{*})=s_{0}.$ Furthermore, without loss of generality, IR and SP allows us to assume that $P^{*}_{i}=P^{s}$.
When $s={\rm{DA}}^{\succ}_{i}(N^{*},P^{*})\,P^{*}_{i}\,\Phi_{i}(N^{*},P^{*})=s_{0}$, WNW implies that $|\Phi_{s}(N^{*},P^{*})|=q_{s}$. Let $\Phi_{s}(N^{*},P^{*})=\{i_{1},\ldots,i_{q_{s}}\}$, where $i_{1}\succ_{s}\cdots\succ_{s}i_{q_{s}}$. Since IR and SP ensure that $\Phi_{i_{1}}(N^{*},(P^{s},P^{*}_{-i_{1}}))=s$, it follows from WLNB that $\Phi_{s}(N^{*},(P^{s},P^{*}_{-i_{1}}))=\Phi_{s}(N^{*},P^{*})$. Applying this argument to $i_{2},\ldots,i_{q_{s}}$, we conclude that there is no loss of generality in assuming that $P^{*}_{j}=P^{s}$ for all $j\in\Phi_{s}(N^{*},P^{*})$. [^19] To adapt Ehlers and Klaus (2016, Theorem 1) to many-to-one matching problems, it is crucial to show that w.l.o.g. it can be assumed that $P^{*}_{j}=P^{s}$ for all $j\in\Phi_{s}(N^{*},P^{*})$ whenever $\Phi_{s}(N^{*},P^{*})\neq{\rm{DA}}^{\succ}_{s}(N^{*},P^{*})$. The same arguments of these authors allow us to show that this property holds for each $j\in\Phi_{s}(N^{*},P^{*})\setminus{\rm{DA}}^{\succ}_{s}(N^{*},P^{*})$ when $\Phi$ satisfies IR, WNW, PM, and SP. Hence, the weak local non-bossiness of $\Phi$ allows us to ensure this property for the students in $\Phi_{s}(N^{*},P^{*})\cap{\rm{DA}}^{\succ}_{s}(N^{*},P^{*})$, if any. Notice that, when $q_{s}=1$, $\Phi_{s}(N^{*},P^{*})\neq{\rm{DA}}^{\succ}_{s}(N^{*},P^{*})$ implies that $\Phi_{s}(N^{*},P^{*})\cap{\rm{DA}}^{\succ}_{s}(N^{*},P^{*})$ is an empty set. Let $N=\{i_{1},\ldots,i_{q_{s}},i\}$. Since $P^{*}_{i}=P^{s}$, $N\subseteq N^{*}$, and ${\rm{DA}}^{\succ}_{i}(N^{*},P^{*})=s$, it follows from PM that ${\rm{DA}}^{\succ}_{i}(N,P^{*})=s$. Moreover, the fact that $P^{*}_{N}=(P^{s},\ldots,P^{s})$ and the definition of ${\rm{DA}}^{\succ}$ guarantee that ${\rm{DA}}^{\succ}_{i_{q_{s}}}(N,P^{*})=s_{0}$ and $i\succ_{s}i_{q_{s}}$. On the other hand, since $N\subseteq N^{*}$ and $\Phi_{s}(N^{*},P^{*})=N\setminus\{i\}$, PM implies that $\Phi_{s}(N,P^{*})=N\setminus\{i\}$. [^20] Remember that $P^{*}_{N}=(P^{s},\ldots,P^{s})=\overline{P}_{N}$ and $\Phi(N,P^{*})$ is independent of the preferences $P^{*}_{\overline{N}\setminus N}$. Hence, $\Phi_{i}(N,P^{*})=s_{0}$ and the property (A) guarantees that $i_{q_{s}}\succ_{s}i$. A contradiction.
When $s=\Phi_{i}(N^{*},P^{*})\,P^{*}_{i}\,{\rm{DA}}^{\succ}_{i}(N^{*},P^{*})=s_{0},$ we can obtain a contradiction applying the arguments of the previous paragraph but swapping the roles of $\Phi$ and ${\rm{DA}}^{\succ}$. $\Box$
Independence of the axioms in Theorem 2. The independence of individual rationality, weak non-wastefulness, population monotonicity, and strategy-proofness is shown in Ehlers and Klaus (2016) through several examples. Since these examples satisfy weakly WrARP and weakly local non-bossiness, they also show that there exist mechanisms that satisfy all the axioms of Theorem 2 except for one of the following: individual rationality, weak non-wastefulness, population monotonicity, or strategy-proofness. Hence, to ensure the independence of the axioms in Theorem 2, it suffices to show that there are mechanisms that satisfy all the axioms but either weakly WrARP or weakly local non-bossiness.
@@
All the axioms are satisfied but weakly WrARP.
$\Phi_{s}(\{1,2,3\},\tilde{P})=\{1,3\},\ \ \quad\quad\Phi_{s}(\{1,2,4\},\tilde{P})=\{2,4\}.$
Moreover, for other $N$ and $P$, assume that $\Phi_{s}(N,P)$ is computed using the serial dictatorship according to the order $(1,2,3,4)$. [^21] The serial dictatorship mechanism follows a given order of agents: the first agent selects her favorite good, the next agent chooses her favorite among the remaining objects, and so on. $\Box$
Example 9.
All the axioms are satisfied but weakly local non-bossiness.
Let $\bar{N}=\{1,2,3\}$, $S=\{s_{1},s_{2}\}$, $q_{s_{1}}=2$, and $q_{s_{2}}=1$.
Consider the profiles $P^{\prime}=(P^{\prime}_{1},P^{\prime}_{2},P^{\prime}_{3})$ and $P^{\prime\prime}=(P^{\prime\prime}_{1},P^{\prime}_{2},P^{\prime}_{3})$, where
$P^{\prime}_{1}:s_{1},s_{2},\quad\quad\quad P^{\prime\prime}_{1}:s_{2},s_{1},\quad\quad\quad P^{\prime}_{2}:s_{1},s_{2},\quad\quad\quad P^{\prime}_{3}:s_{1},s_{2}.$
Let $\Phi$ be the mechanism such that
$\Phi_{s_{1}}(\{1,2,3\},P^{\prime})=\{1,2\},\ \quad\quad\quad\Phi_{s_{2}}(\{1,2,3\},P^{\prime})=\{3\},$
$\Phi_{s_{1}}(\{1,2,3\},P^{\prime\prime})=\{1,3\},\,\,\,\quad\quad\quad\Phi_{s_{2}}(\{1,2,3\},P^{\prime\prime})=\{2\}.$
Moreover, for other $N$ and $P$, assume that $\Phi(N,P)$ is computed using the serial dictatorship according to the order $(1,2,3)$. $\Box$

### v5

@@ Appendix B On local non-bossiness under non-responsive preferences
We claim that the mechanism that associates to each preference profile the school-median stable matching is locally bossy. Let $N=\{1,2,3,4\}$, $S=\{s_{1},s_{2}\}$, $q_{s_{1}}=q_{s_{2}}=2$, $\succ_{s_{1}}:2,1,3,4$, and $\succ_{s_{2}}:3,4,2,1.$ If preferences are given by $P_{1}:s_{2},s_{1},s_{0}$, $P_{2}:s_{2},s_{1},s_{0}$, $P_{3}:s_{1},s_{2},s_{0}$, and $P_{4}:s_{1},s_{2},s_{0}$, then there are three stable matchings:
@@
It follows that $\eta$ is the school-median stable matching. Now, when student $1$’s preferences change to $P^{\prime}_{1}:s_{1},s_{2},s_{0}$, the only stable matchings are $\eta$ and $\rho$. In this case, $\rho$ is the school-median stable matching. Therefore, as $\eta(1)=\rho(1)=s_{1}$ and $\eta(s_{1})\neq\rho(s_{1})$, we conclude that the school-median stable mechanism is locally bossy. $\Box$
The next example shows that the construction of the graph $G^{\prime}$ in the proof of Theorem 1 is not superfluous. Indeed, the $\mu$-improving cycle $(i^{*}_{1},\ldots,i^{*}_{r})$, that allows us to show that $\mu(\overline{s})=\mu^{\prime}(\overline{s}),$ may not be present in the graph $G$.
@@
Let $N=\{1,2,3,4,5,6\}$, $S=\{s_{1},s_{2},s_{3},s_{4},s_{5}\}$, $q_{s_{1}}=2$, $q_{s}=1$ for every $s\neq s_{1}$. and
$\quad\succ_{s_{1}}:6,1,\ldots,\quad\quad\succ_{s_{2}}:5,1,3,2\ldots,\quad\quad\succ_{s_{3}}:4,2,3\ldots\quad\quad\succ_{s_{4}}:3,4\ldots\quad\quad\succ_{s_{5}}:2,5\ldots$
Consider a preference profile $P\in{\mathcal{P}}$ such that
$\begin{array}[]{ll}P_{1}:&s_{2},s_{1},\ldots,\\ P_{2}:&s_{2},s_{3},s_{5},\ldots,\end{array}$ $\quad\quad\quad\begin{array}[]{ll}P_{3}:&s_{3},s_{2},s_{4},\ldots,\\ P_{4}:&s_{4},s_{3},\ldots,\end{array}$ $\quad\quad\quad\begin{array}[]{ll}P_{5}:&s_{5},s_{2},\ldots,\\ P_{6}:&s_{1},\ldots,\end{array}$
and let $P^{\prime}_{1}:s_{1},\ldots$. Then
$\mu$ $\equiv$ ${\rm{DA}}(P)=((1,s_{1}),(2,s_{5}),(3,s_{4}),(4,s_{3}),(5,s_{2}),(6,s_{1})),$
$\mu^{\prime}$ $\equiv$ ${\rm{DA}}(P^{\prime}_{1},P_{-1})=((1,s_{1}),(2,s_{2}),(3,s_{3}),(4,s_{4}),(5,s_{5}),(6,s_{1})).$
Using the notation from the proof of Theorem 1, it follows that $G=(V,E)$ is defined by $V=\{2,3,4,5\}$ and $E=\{[2,5],[3,4],[4,3],[5,2]\}$. Moreover, the graph $G^{\prime}=(V,E^{\prime})$ satisfies $E^{\prime}=\{[3,5],[2,4],[4,3],[5,2]\}.$ Hence, $G^{\prime}$ has a unique $\mu$-improving cycle, namely: $(2,4,3,5)$. As expected, student $1$ is the only student who $\mu$-blocks this cycle (at edge $[3,5]$). On the other hand, $G$ has two $\mu$-improving cycles: $(2,5)$ and $(3,4)$. The first one is $\mu$-blocked by the students 11 and $3$ at edge $[2,5]$, while the second one is $\mu$-blocked by the student $2$ at edge $[3,4]$. More importantly, $\mu^{\prime}$ does not implement $(2,4,3,5)$. $\Box$
In 11, student $1$ has a relevant role in the inefficiency of ${\rm{DA}}(P)$. Indeed, although $1$ is assigned to $s_{1}$, she displaces $2$ from $s_{2}$. However, the displacement of $2$ induces other inefficiencies: although $2$ is assigned to $s_{5}$, she displaces $3$ from $s_{3}$. Hence, when student $1$ reports $P^{\prime}_{1}$, the improvement in welfare with respect to ${\rm{DA}}(P)$ is greater than the one generated when only $(2,4,3,5)$ is implemented. Indeed, the matching obtained from ${\rm{DA}}(P)$ when $(2,4,3,5)$ in implemented, $\eta=((1,s_{1}),(2,s_{3}),(3,s_{2}),(4,s_{4}),(5,s_{5}),(6,s_{1}))$, corrects the inefficiencies generated after the displacement of $3$ from $s_{3}$. Therefore, in the context of the example above, ${\rm{DA}}(P^{\prime}_{1},P_{-1})$ Pareto dominates the matching obtained when a $\mu$-improving cycle present in $G^{\prime}$ but not in $G$ is implemented from ${\rm{DA}}(P)$.
This last property always holds. More formally, let $(i^{*}_{1},\ldots,i^{*}_{r})$ be a $\mu$-improving cycle present in $G^{\prime}$ (see the proof of Theorem 1, Case I). We affirm that $\mu^{\prime}$ weakly Pareto dominates under $P$ the matching $\eta$ obtained by implementing $(i^{*}_{1},\ldots,i^{*}_{r})$ from $\mu$. By contradiction, assume that $\eta(k)P_{k}\mu^{\prime}(k)$ for some $k\in N$. Since $\mu^{\prime}$ Pareto dominates $\mu$ under $P$, it follows that $k=i^{*}_{s}$ for some $s\in\{1,\ldots,r\}$, $k$ $\mu$-blocks an edge $i,i^{*}_{s+1}]$ of $G$, and $\eta(k)=\mu(i^{*}_{s+1})=\mu^{\prime}(i)$ [modulo $r$]. Hence, if $s=\mu^{\prime}(i)$, then $k\succ_{s}i$ and $sP_{k}\mu^{\prime}(k)$. This contradicts the stability of $\mu^{\prime}$.
Kesten (2010) introduces the efficiency adjusted deferred acceptance mechanism ($\rm{EADAM}$}). This mechanism implements improvement cycles sequentially in such a way that $\rm{DA}$} is recalculated after modifying the preferences of some students who agree to waive their priorities at schools. In particular, when everyone agrees to waive their priorities at schools, $\rm{EADAM}$} is efficient and Pareto dominates $\rm{DA}$}.
Our results are not directly related to Kesten (2010). For instance, the improvement cycle constructed in the proof of Theorem 1—the one that is blocked only by the student who misreports her preferences—does not necessarily coincide with one of the improvement cycles implemented by $\rm{EADAM}$}. Indeed, it is not difficult to verify that in 11 we have that
${\rm{EADAM}}(P)=((1,s_{1}),(2,s_{2}),(3,s_{3}),(4,s_{4}),(5,s_{5}),(6,s_{1})).$
Hence, ${\rm{EADAM}}(P)$ does not implement the $\mu$-improving cycle $(2,4,3,5)$.